# Distinguishing Calabi-Yau Topology using Machine Learning

### v1

@@ 3.2 Architectures
Our model is displayed in the Figure 1, which takes the configuration matrix of the CICY manifolds as input and outputs a 1$\times$4 vector, representing the predicted CICY triple intersection numbers $d_{1}$, $d_{2}$, $d_{3}$ and $d_{p}$. The architecture consists of three inception modules with 32, 64, and 32 filters, respectively. Within each inception module, we employ the rows 12$\times$1 and the columns 1$\times$15 kernels for two parallel convolutions with a stride of 1 and concatenate the outputs of both layers together over the channel dimension. The Leaky ReLU function with negative slope 0.2 serves as the activation function.
To maintain consistent data dimensions, we attempt zero-padding and adopt the same padding options. Batch normalization layers follow each concatenation layer, facilitating the connection between these inception modules. The output from the final inception module is passed through a flatten layer, which then connects to a fully connected layer that outputs a four-dimensional vector. During training, we utilize the Adam gradient descent optimizer with Mean Square Error (MSE) as the loss function. The initial learning rate is set at 0.0001, and the batch size was specified as 32. To mitigate overfitting, we introduce $l_{2}$ regularization with weight 0.002. This comprehensive approach aims to enhance the model’s predictive accuracy while addressing potential overfitting challenges. The model was trained over 3000 epochs to ensure thorough learning and optimization of parameters.
@@ 4 Results
Figure 4: This presents a comparative analysis between our Inception Network model and conventional machine learning regressors across various fraction of training data. The first figure presents a line graph measuring the accuracy of different regressors in different fractions of training data while the second figure presents MSE.
@@
Moreover, we conduct a comprehensive comparison between our deep learning model and several classic regression models. This comparison is aimed at evaluating the relative effectiveness and accuracy of our model in contrast to traditional methods. The result demonstrate that deep learning methods significantly out-perform other traditional regression methods as shown in Figure 3. Furthermore, we delve into an extensive investigation of our model’s performance across various ratios of training set splits. This exploration provides insights into the adaptability and robustness of our model under different training conditions. We find that our model is capable of making good predictions across various training set splits compared with other regressors and display the results in Figure 4.
@@
Table 1 provides a comprehensive analysis of our model’s performance on the individual intersection numbers $d_{1}$, $d_{2}$, $d_{3}$, and $d_{p}$. It highlights that our model achieves high accuracy for each of these CICY triple intersection numbers. Additionally, the MSE across these dimensions is kept low, indicating that the model not only predicts with high precision but also with minimal error, thus demonstrating robust performance even in the handling of complex manifold structures. This effective combination of high accuracy and low MSE underscores the efficacy of our model in dealing with the intricacies of CICY manifold predictions. We can observe that $d_{p}$ exhibits relatively a bit lower prediction accuracy and higher MSE, which stems from the more intricate structure of $d_{p}$ possibly.
In Figure 3, we employed some distinct classic machine learning regression models to predict the CICY triple intersection numbers $d_{1}$, $d_{2}$, $d_{3}$ and $d_{p}$, and compared their performance with our model. These algorithms include SVM regressor, Random Forest regressor and XGBoost regressor, which served as benchmark models. The settings of these models are as follows.
@@ 5 Discussion
We recommend future research to further refine the model’s predictive capabilities, explore additional properties of CICY manifolds, and extend the application of deep learning models to a wider array of mathematical structures.

### v2

@@ 3.2 Architectures
[…]
To maintain consistent data dimensions, we attempt zero-padding and adopt the same padding options. Batch normalization layers follow each concatenation layer, facilitating the connection between these inception modules. The output from the final inception module is passed through a flatten layer, which then connects to a fully connected layer that outputs a four-dimensional vector. During training, we utilize the Adam gradient descent optimizer with Mean Square Error (MSE) as the loss function. The initial learning rate is set at 0.0001, and the batch size was specified as 32. To mitigate overfitting, we introduce $l_{2}$ regularization with weight 0.002. This comprehensive approach aims to enhance the model’s predictive accuracy while addressing potential overfitting challenges. The model was trained over 3000 epochs to ensure thorough learning and stable optimization of the parameters, during which we observed consistent improvements in performance metrics until convergence.
@@ 4 Results
Figure 3: This residual scatter plots comparing true and predicted CICY triple intersection numbers $d_{1}$, $d_{2}$, $d_{3}$, and $d_{p}$, based on an 80%/20% train–test split.
@@
Figure 5: This presents a comparative analysis between our Inception Network model and conventional machine learning regressors across various fractions of training data. The first figure presents a line graph measuring the accuracy of different regressors in different fractions of training data while the second figure presents MSE.
@@
Moreover, we conduct a comprehensive comparison between our deep learning model and several classic regression models. This comparison is aimed at evaluating the relative effectiveness and accuracy of our model in contrast to traditional methods. The results demonstrate that deep learning methods significantly out-perform other traditional regression methods as shown in Figure 4. Furthermore, we delve into an extensive investigation of our model’s performance across various ratios of training set splits. This exploration provides insights into the adaptability and robustness of our model under different training conditions. We find that our model is capable of making good predictions across various training set splits compared with other regressors and display the results in Figure 5.
@@
Figure 3 presents residual scatter plots comparing the true and predicted values of the CICY triple intersection numbers $d_{1}$, $d_{2}$, $d_{3}$, and $d_{p}$. Each subplot displays the deviation $(\text{True}-\text{Predicted})$ with respect to the true values, providing a sample-wise view of model accuracy. The residuals are generally centered around zero, indicating that the model captures the underlying distribution of the data well. Slightly larger deviations are observed for $d_{3}$ and $d_{p}$, suggesting that these quantities exhibit higher structural complexity and thus pose greater learning difficulty for the network.
Table 1 provides a comprehensive analysis of our model’s performance on the individual intersection numbers $d_{1}$, $d_{2}$, $d_{3}$, and $d_{p}$. It highlights that our model achieves high accuracy for each of these CICY triple intersection numbers. Additionally, the MSE across these dimensions is kept low, indicating that the model not only predicts with high accuracy but also with minimal error, thus demonstrating robust performance even in the handling of complex manifold structures. This effective combination of high accuracy and low MSE underscores the efficacy of our model in dealing with the intricacies of CICY manifold predictions. We can observe that $d_{p}$ exhibits relatively a bit lower prediction accuracy and higher MSE, which stems from the more intricate structure of $d_{p}$ possibly. In particular, $d_{1}$ is almost always $1$ or $2$, making its distribution highly concentrated and relatively easy to predict. By contrast, $d_{p}$ is derived from contractions of the second Chern class and spans a much broader range of values, up to $64$.
In Figure 4, we employed some distinct classic machine learning regression models to predict the CICY triple intersection numbers $d_{1}$, $d_{2}$, $d_{3}$ and $d_{p}$, and compared their performance with our model. These algorithms include SVM regressor (cortes1995support), Random Forest regressor (breiman2001random) and XGBoost regressor (chen2016xgboost), which served as benchmark models. The settings of these models are as follows.
@@
We note that for all baseline regressors (SVM, Random Forest, XGBoost), we did not perform extensive hyperparameter tuning beyond standard package defaults or common settings. While this provides a fair benchmark comparison, their performance could likely be further improved with systematic hyperparameter optimization.
@@ 5 Discussion
We recommend future research to further refine the model’s predictive capabilities including utilizing the recent developed manifold fitting techniques (yao2019manifold; yao2023manifold; yao2023pnas; yao2024pnas), explore additional properties of CICY manifolds, and extend the application of deep learning models to a wider array of mathematical structures. For instance, a nice work (coates2023machine) explored the quantum dimension of Fano varieties, it would be a very fruitful study to be able to predict more refined objects such as Gromow-Witten invariants of our CICY manifolds.
The ultimate goal of this study, as with all AI-driven mathematical discovery (He:2019nzx; He:2024gnk), is to uncover new, interpretable structures, either in the form of a conjectured formula or a new derivation/proof. A clear future work is to extract a conjectural formula amongst $d_{1,2,3,p}$ in an automated way, that would pass the Automaticity, Interpretability, and Non-triviality of the Birch test of AI-driven findings (he2024can).